OpenStudy (anonymous):

Simplify the expression. State any restrictions on the variable. x^2+6x+8/x^2-10x-24

OpenStudy (anonymous):

x^2 +6x +8 = (x+2)(x+4) x^2-10x-24 = (x+2)(x-12) x^2+6x+8/x^2-10x-24 =(x+2)(x+4) / (x+2)(x-12) =(x+4)/(x-12) (x-12)=0 x=12 denominator cannot be zero so x cannot be 12